OpenStudy (anonymous):

x^3-729 factor the difference of the two cubes

OpenStudy (amistre64):

(a^3 - b^3) (a-b)(a^2-ab+b^2) ?

OpenStudy (anonymous):

i am lost

OpenStudy (amistre64):

what is the cuberoot of 729?

OpenStudy (anonymous):

ummm

OpenStudy (amistre64):

9....got it; (x-9)(x^2 +9x +81) or is it... (x-9)(x^2 -9x +81) one or the other, I forget which :)

myininaya (myininaya):

its the first one

OpenStudy (amistre64):

hey... i just figured that out :)

myininaya (myininaya):

a^3+b^3=(a+b)(a^2-ab+b^2) and a^3-b^3=(a-b)(a^2+ab+b^2)

OpenStudy (amistre64):

yep.... learned it like 4 semesters ago and havent used it since :/

myininaya (myininaya):

i always remember that is one or the other, but then i think if i set this =0 and solve for x, I would get x=9 so one of the factors is x-9 since x-9=0